A manager believes that the demand for his product is given by the equation P=50-Q/100.
a. what is the elasticity of demand as price decreases from price12 to price 10?
B. what is the elasticity of demand as price increases from price10 to price 12?
1
Expert's answer
2020-04-10T09:13:26-0400

a. If price decreases from p1 = 12 to p2 = 10, then Q1 = 5,000 - 100×12 = 3,800 units, Q2 = 5,000 - 100×10 = 4,000 units.

"Ed = \frac {4,000 - 3,800}{10 - 12}×\frac{12+10}{3,800 + 4,000} = -0.28."

So, the demand is inelastic.

B. If price increases from p1 = 10 to p2 = 12, then Q1 = 5,000 - 100×10 = 4,000 units, Q2 = 5,000 - 100×12 = 3,800 units.

"Ed = \frac {3,800 - 4,000}{12 - 10}×\frac{12+10}{3,800 + 4,000} = -0.28."

So, the demand is inelastic.